OpenStudy (anonymous):

in 90 miles in 75 minutes how far can they travel in 2 hour

OpenStudy (anonymous):

135

OpenStudy (anonymous):

75+75=150 which is a hour and a half so 90*2=180 then -45=135

OpenStudy (anonymous):

Do porportions\[\frac{ 75\min }{ 90miles }=\frac{ 120\min }{ xmiles }=>10800=75x=>\frac{ 10800 }{ 75 }=x\]